OpenStudy (anonymous):

The area of the base of a prism is 70 mm2. The perimeter of the base is 50 mm. The height of the prism is 9 mm. What is the surface area of the prism? A. 129 mm2 B. 520 mm2 C. 590 mm2 D. 730 mm2

OpenStudy (anonymous):

surface area of a rectangular prism with a base length of 70 mm^2, perimeter of the base is 50 mm. The height of the prism is 9 mm. let a= base length, b = base width, c = height=9. a*b= 70. a+b+9=50, or a+b=41, now bc+ca= c(a+b)=9*41=369 so suface area= 2(ab+bc+ca)= 2(70+369)=878 mm^2